OpenStudy (anonymous):

40 books in 16 minutes what is the unit rate

OpenStudy (unklerhaukus):

the unit rate is the number of books per 1 minute

OpenStudy (anonymous):

which is a better buy $48 that has 15 pounds or $50 that has 16 pounds and why

OpenStudy (unklerhaukus):

what is the unit value of one pound in each

OpenStudy (anonymous):

15 pounds comes out to 3.2 and 16 pounds comes out to 3.125

OpenStudy (unklerhaukus):

good, so the better buy is the one that has better value, ie less cost for the same amount

OpenStudy (anonymous):

thanks lola has 7 shirts but only takes 2 with her what is the % of shirts she takes with her

OpenStudy (anonymous):

4 over 100 is 40% right or 4%

OpenStudy (unklerhaukus):

\[\frac 27=\frac{2}{7}\times\frac{100}{100}=\frac{200}{700}=\frac{200/7}{100}\] \[=\frac{210/7-10/7}{100}=\frac{30-1-3/7}{100}=\frac{29}{100}-\frac {3/7}{100}\approx29\%\]

OpenStudy (unklerhaukus):

\[\frac{4}{100}=4\% \]

OpenStudy (anonymous):

thanks a million what is the estimate result of 59 5/7 divided by 4 7/12

OpenStudy (anonymous):

.52 X 43 is the same as 52% of 43?